# IntactKV: Improving Large Language Model Quantization by Keeping Pivot Tokens Intact

### v1

@@ Abstract
Large language models (LLMs) excel in natural language processing but demand intensive computation. To mitigate this, various quantization methods have been explored, yet they compromise LLM performance. This paper unveils a previously overlooked type of outlier in LLMs. Such outliers are found to allocate most of the attention scores on initial tokens of input, termed as pivot tokens, which is crucial to the performance of quantized LLMs. Given that, we propose IntactKV to generate the KV cache of pivot tokens losslessly from the full-precision model. The approach is simple and easy to combine with existing quantization solutions. Besides, IntactKV can be calibrated as additional LLM parameters to boost the quantized LLMs further. Mathematical analysis also proves that IntactKV effectively reduces the upper bound of quantization error. Empirical results show that IntactKV brings consistent improvement and achieves lossless weight-only INT4 quantization on various downstream tasks, leading to the new state-of-the-art for LLM quantization.
@@
[…]
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in various tasks and benchmarks in natural language processing Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023a); Team et al. (2023). Nonetheless, the rise of LLMs also increases computational intensity and memory requirements. This motivates various research to decrease the inference cost of LLMs, e.g., quantizaiton Frantar et al. (2022); Shao et al. (2024); Lin et al. (2023), pruning Frantar and Alistarh (2023); Liu et al. (2023b); Sun et al. (2023), and speculative decoding Chen et al. (2023); Leviathan et al. (2023); Cai et al. (2024), e.t.c.
@@
Towards that end, we are motivated to propose IntactKV, a simple strategy that is orthogonal to most existing quantization solutions. The key idea behind IntactKV is to generate the KV cache of pivot tokens from the full-precision model. By keeping the KV cache of pivot tokens intact, quantization error accumulated on the output of self-attention will be effectively alleviated in the rest of the decoding steps. Moreover, IntactKV can also serve as extra trainable parameters in addition to the LLM backbone. The calibration process of IntactKV follows the convention of PTQ Bai et al. (2022); Frantar et al. (2022); Lin et al. (2023), which further decreases the quantization error. To get more insights from IntactKV, we also provide mathematical analysis and the results show that IntactKV can effectively lower the upper bound of quantization error.
Empirical results show that IntactKV consistently improves the capability of quantized models on various open-sourced LLMs (e.g., LLaMA and Vicuna) and different downstream tasks such as C4, MMLU, commonsense QA, and MT-bench. When armed with AWQ Lin et al. (2023), IntactKV achieves new state-of-the-art quantization results, e.g., lossless INT4 weight-only quantization for Vicuna-v1.5 on commonsense QA tasks. Moreover, fine-tuning IntactKV with INT4 quantization even matches the full-precision model on aligning with human preference, as evaluated by GPT-4 Bubeck et al. (2023) on MT-bench.
@@ 2.1 Preliminaries on LLM Quantization
Figure 3: The overview of the proposed IntactKV applied for the supervised fine-tuned LLM. The full-precision model takes the system prompt as input and generates the IntactKV losslessly as the prefix concatenated with the rest of the KV cache of quantized LLMs. IntactKV can be further calibrated by minimizing the mean square error $\mathcal{L}$ between the full-precision and quantized LLM.
@@ 3.1 Preserving the KV Cache of Pivot Tokens
According to Section 2.2, the attention sinks of pivot tokens are likely to deteriorate by quantization. To alleviate this issue, we propose IntactKV, a simple and effective strategy to keep these pivot tokens intact. Specifically, as illustrated in Figure 2(a), we leverage the full-precision LLM to generate the lossless KV cache of pivot tokens, which is saved offline. The quantized LLM then loads IntactKV as the prefix to concatenate with the rest of the KV cache and continues with the regular auto-regressive decoding. The pseudo code of the inference scheme with IntactKV is presented in Figure 2(b).
In order to study the benefits of IntactKV, we conduct a preliminary test on the mean square error (MSE) of the attention and transformer layer output. From Figure 2, it is natural that the increasing size of IntactKV gives the monotonically decreasing MSE on both the attention and transformer layers. More importantly, it is found the same tokens in Section 2.2 (e.g., [BOS] and other delimiter tokens) give the most significant decrease on the MSE, which demonstrates the importance of their KV cache. This aligns with the observations in Figure 1 that pivot tokens exhibit outliers with extreme values and attention sinks.
@@ 3.2 IntactKV as Trainable Parameters
Since IntactKV is pre-computed and saved offline, it can be treated as extra trainable parameters aside from the LLM backbone to further boost the quantized LLMs. Despite there being no information loss at the pivot tokens, the quantization may still introduce errors to the KV cache during the decoding stage. As shown in Figure 2(a), we calibrate IntactKV to compensate for the quantization error accumulated in the following tokens. While there are various metrics to characterize the quantization discrepancy Frantar et al. (2022); Shao et al. (2024); Liu et al. (2023a), we adopt the mean square error of the transformer layer output between the full-precision LLM and quantized LLM, a simple yet most widely used metric, i.e.,
@@
where $\Theta$ denotes the set of IntactKV, $f_{l}$ is the mapping function for the $l$-th Transformer layer, and $L$ is the number of Transformer layers in LLM. $\mathbf{x}$ is the input sequence, while $\mathbf{w},\hat{\mathbf{w}}$ are full-precision and quantized weights respectively. Note that the full-precision model is only required during the calibration process, and it can be then discarded afterward. It is empirically found that calibration of system prompt IntactKV in SFT models generally gives more improvement than the calibration of [BOS] token in pre-trained LLMs. This matches the intuition that a larger size of IntactKV increases the potential to compensate for quantization errors.
As we focus on the post-training quantization, the training of IntactKV is highly lightweight since the only learnable parameters introduced are IntactKV, i.e., the KV cache of pivot tokens. It takes only as few as 20 epochs on a calibration set with 128 samples. Besides, training with a quantized model further lowers the memory cost.
@@ 3.3 Theoretical Analysis
The proof to Theorem 1 can be found in Appendix A. We preserve the terms w.r.t. $\Delta{\bm{K}}$ and $\Delta{\bm{V}}$ of interests, and leave the rest as constants. Note that $\Delta{\bm{K}}$ can be further separated by the pivot tokens $\Delta{\bm{K}}_{p}$ and rest tokens $\𝑝\Delta{\bm{K}}_{\textbackslash p}$, and similar notations hold for $\Delta{\bm{V}}$. Therefore, we have ${\|\Delta{\bm{K}}\|}_{2,\infty}=\max\big{(}{\|\Delta{\bm{K}}_{p}\|}_{2,\infty},{\|\Delta{\bm{K}}_{\textbackslash p}\|}_{2,\infty}\big{)},$ and ${\|\Delta{\bm{V}}\|}_{F}=\sqrt{{\|\Delta{\bm{V}}_{p}\|}_{F}^{2}+{\|\Delta{\bm{V}}_{\textbackslash p}\|}_{F}^{2}}$. With IntactKV  we have ${\|\Delta{\bm{K}}_{p}\|}_{2,\infty}=\|\Delta{\bm{V}}_{p}\|_{F}=0$ since they are generated losslessly, which decreases the upper bound of $\|\Delta{\bm{h}}\|_{2}$. Moreover, it can further reduce the bound by incorporating more pivot tokens. This also aligns with the observation in Figure 2 that a larger size of IntactKV gives a lower MSE of the attention module.
@@ 4 Experiments
FP16 7.36 6.82 6.15 5.83 7.11 6.58 5.59
RTN 9.15 7.89 6.85 6.33 8.79 7.43 6.12
GPTQ 8.59 7.49 6.73 6.29 9.41 7.25 6.27
OmniQuant 8.24 7.38 6.64 6.17 8.21 7.23 5.96
AWQ 8.26 7.38 6.59 6.16 8.12 7.15 5.91
  +IntactKV${}_{\textrm{[B]}}$ 8.12 7.36 6.54 6.12 8.00 7.12 5.89
Table 1: INT3-group128 quantization results of LLaMA and LLaMA-2 Models on C4 dataset.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
OmniQuant 49.52% 41.22% 59.47% 57.74% 51.82% 49.12% 44.40% 60.48% 58.95% 52.86%
Table 2: INT3-group128 quantization results of Vicuna-v1.5-13B on 0-shot and 5-shot MMLU benchmarks.
@@
RTN 44.62% 51.44% 39.33% 44.56% 53.18% 60.99% 65.40% 59.13% 63.23% 67.19%
GPTQ 43.99% 52.95% 40.12% 47.83% 55.84% 58.09% 65.78% 59.80% 64.03% 66.68%
OmniQuant 46.54% 52.86% 43.18% 47.92% 55.12% 61.73% 65.23% 61.08% 64.81% 67.58%
AWQ 46.45% 52.92% 43.08% 48.56% 56.09% 61.86% 66.01% 60.94% 64.53% 67.67%
@@
Table 3: INT3-group128 quantization results of various Vicuna models on 5-shot MMLU and 0-shot QA tasks.
@@ 4.1 Settings
We evaluate the proposed IntactKV on various sizes of open-sourced LLMs, including LLaMA (Touvron et al., 2023a) (7B-65B), LLaMA-2 (Touvron et al., 2023b) (7B-70B), Vicuna-v1.3 (7B-33B) and Vicuna-v1.5 (Touvron et al., 2023b) (7B-13B). We denote models that keep intact [BOS] KV as IntactKV${}_{\textrm{[B]}}$, and models that keep intact system prompt KV as IntactKV${}_{\textrm{[P]}}$.
@@
We mainly consider weight-only quantization methods, including round-to-nearest quantization (RTN), GPTQ (Frantar et al., 2022), the state-of-the-art OmniQuant (Shao et al., 2024) and AWQ (Lin et al., 2023). For GPTQ, we use AutoGPTQ [^1] https://github.com/AutoGPTQ/AutoGPTQ with C4 calibration set following GPTQ paper (Frantar et al., 2022) to reproduce all results. For AWQ [^2] https://github.com/mit-han-lab/llm-awq and OmniQuant [^3] https://github.com/OpenGVLab/OmniQuant, we directly load the officially released quantization parameters of LLaMA models for evaluation and reproduce results on Vicuna models with their official code. We use Pile (Gao et al., 2020) calibration set for AWQ and WikiText2 (Merity et al., 2016) calibration set for OmniQuant, following (Lin et al., 2023; Shao et al., 2024). We adopt asymmetric group-wise quantization with a group size of 128 and mainly focus on INT3 and INT4 quantization since INT8 is empirically lossless on various task metrics (Dettmers et al., 2022).
While IntactKV can be readily combined with these existing quantization methods, we mainly apply IntactKV for AWQ in the main experiments due to the space limitation and the state-of-the-art performance of AWQ. We shall provide a more comprehensive evaluation of combining IntactKV with other quantization methods in Section 4.3. Moreover, aside from weight-only quantization, the proposed IntactKV can be similarly applied for KV cache quantization, as detailed in Section 4.4. Nonetheless, it is non-trivial to apply IntactKV for activation quantization, since the quantization over KV cache cannot keep them intact. We leave it as future work and more discussions are provided in Section 6.
@@
For pre-trained LLMs (i.e., LLaMA and LLaMA-2), we report the perplexity (PPL) of language generation on C4 (Raffel et al., 2020) and WikiText2 (Merity et al., 2016) dataset. For SFT models (i.e., Vicuna-v1.3 and v1.5), we conduct evaluation over a wide range of downstream tasks. We test the zero and five-shot performance on the Massively Multitask Language Understanding (MMLU) (Hendrycks et al., 2020) benchmark. Meanwhile, we also evaluate seven zero-shot commonsense QA tasks: OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), ARC-Challenge, ARC-Easy (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), and LAMBADA (Paperno et al., 2016). Additionally, we evaluate quantized Vicuna on MT-bench (Zheng et al., 2023), a high-quality dataset consisting of 80 open-ended multi-turn questions, to gauge their alignment with human preferences. The responses generated by quantized models are judged by GPT-4 with a total score of 10. More evaluation details can be found in Appendix E.
@@
w3g128 RTN 42.00% 70.01% 47.44% 72.77% 82.20% 74.18% 69.20% 65.40%
GPTQ 42.40% 69.53% 48.46% 71.84% 83.76% 74.48% 70.00% 65.78%
OmniQuant 43.40% 68.51% 47.53% 71.09% 82.32% 73.92% 69.84% 65.23%
AWQ 44.00% 68.75% 48.21% 71.76% 83.58% 75.09% 70.68% 66.01%
w4g128 RTN 44.20% 70.80% 48.98% 73.82% 84.68% 76.36% 73.04% 67.41%
GPTQ 45.80% 70.96% 50.51% 73.99% 85.47% 76.70% 73.43% 68.12%
OmniQuant 43.80% 70.24% 49.74% 73.61% 84.59% 76.35% 72.54% 67.27%
AWQ 44.00% 72.06% 49.15% 73.44% 85.17% 77.00% 72.77% 67.66%
Table 4: Quantization results of Vicuna-v1.5-13B on seven 0-shot commonsense QA tasks.
For evaluation on PPL, MMLU, and commonsense QA tasks, we adopt IntactKV${}_{\textrm{[B]}}$ that only includes [BOS] KV since the input sequence of these tasks does not use any system prompt. For evaluation of SFT models on MT-bench, we adopt IntactKV${}_{\textrm{[P]}}$ to keep an intact system prompt KV cache. The system prompt of Vicuna can be found in Appendix B. For training the cached IntactKV, we randomly sample 128 samples from ShareGPT [^4] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered dataset as our calibration dataset, consisting of multi-turn ChatGPT (OpenAI, 2022) conversations. The layer-wise MSE defined in Equation 2 is calculated on the response of ChatGPT. We use AdamW optimizer with learning rate $2\times 10^{-4}$, training for 160 optimizer update steps with a gradient accumulation step of 16, i.e., 20 epochs. As mentioned in Section 3.2, training IntactKV${}_{\textrm{[B]}}$ leads to comparable performance compared with vanilla IntactKV. Instead, the calibration of IntactKV${}_{\textrm{[P]}}$ has more potential to improve quantized LLMs with longer system prompt. Thus, we primarily evaluate the trainable IntactKV${}_{\textrm{[P]}}$ with system prompt as pivot tokens in the following experiments.

### v2

@@ Abstract
Large language models (LLMs) excel in natural language processing but demand intensive computation. To mitigate this, various quantization methods have been explored, yet they compromise LLM performance. This paper unveils a previously overlooked type of outliers in LLMs. Such outliers are found to allocate most of the attention scores on initial tokens of input, termed as pivot tokens, which are crucial to the performance of quantized LLMs. Given that, we propose IntactKV to generate the KV cache of pivot tokens losslessly from the full-precision model. The approach is simple and easy to combine with existing quantization solutions with no extra inference overhead. Besides, IntactKV can be calibrated as additional LLM parameters to boost the quantized LLMs further with minimal training costs. Mathematical analysis also proves that IntactKV effectively reduces the upper bound of quantization error. Empirical results show that IntactKV brings consistent improvement over various quantization methods across different LLMs and downstream tasks, leading to the new state-of-the-art for LLM quantization. The codes are available at https://github.com/ruikangliu/IntactKV.
@@
Ruikang Liu1       Haoli Bai2       Haokun Lin3       Yuening Li4       Han Gao2 Zhengzhuo Xu1       Lu Hou2       Jun Yao2       Chun Yuan1†thanks: † Corresponding author. 1Shenzhen International Graduate School, Tsinghua University    2Huawei Noah’s Ark Lab 3Institute of Automation, Chinese Academy of Sciences    4The Chinese University of Hong Kong {liuruikang.cs, xzzthu}@gmail.com    {baihaoli, han.g, houlu3, yaojun97}@huawei.com haokun.lin@cripac.ia.ac.cn    yuening@link.cuhk.edu.hk    yuanc@sz.tsinghua.edu.cn
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in various tasks and benchmarks in natural language processing Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023a); Team et al. (2023). Nonetheless, the rise of LLMs also increases computational intensity and memory requirements. This motivates various research to decrease the inference cost of LLMs, e.g., quantizaiton Frantar et al. (2022); Shao et al. (2024); Lin et al. (2023), pruning Frantar and Alistarh (2023); Liu et al. (2023b); Sun et al. (2023); Zhang et al. (2024), and speculative decoding Chen et al. (2023); Leviathan et al. (2023); Cai et al. (2024), e.t.c.
@@
Towards that end, we are motivated to propose IntactKV, a simple strategy that is orthogonal to most existing quantization solutions. The key idea behind IntactKV is to generate the lossless KV cache of pivot tokens from the full-precision model. By keeping the KV cache of pivot tokens intact, quantization error accumulated on the output of self-attention will be effectively alleviated in the rest of the decoding steps. The integration of IntactKV comes with no additional inference overhead. Moreover, IntactKV can also serve as extra trainable parameters in addition to the LLM backbone. The calibration process of IntactKV follows the convention of PTQ Bai et al. (2022); Frantar et al. (2022), which further decreases the quantization error. To get more insights from IntactKV, we also provide mathematical analysis and the results show that IntactKV can effectively lower the upper bound of quantization error.
Empirical results show that IntactKV consistently improves the capability of different quantization methods (e.g. AWQ Lin et al. (2023), GPTQ Frantar et al. (2022), OmniQuant Shao et al. (2024) and QuaRot Ashkboos et al. (2024)) on various open-sourced LLMs (e.g., LLaMA and Vicuna) across different tasks and benchmarks such as PPL, MMLU, commonsense QA, and MT-bench, achieving new state-of-the-art results for weight-only quantization as well as weight and activation quantization, e.g., lossless INT4 weight-only quantization for Vicuna-v1.3-13B on commonsense QA tasks. Moreover, calibrating IntactKV with INT4 quantization even matches the full-precision model on aligning with human preferences, as evaluated by GPT-4 Bubeck et al. (2023) on MT-bench.
@@ 2.1 Preliminaries on LLM Quantization
Figure 3: The overview of the proposed IntactKV applied for the supervised fine-tuned LLM. The full-precision model takes the system prompt as input and generates the IntactKV losslessly as the prefix concatenated with the rest of the KV cache of quantized LLMs. IntactKV can be further calibrated by minimizing the mean squared error $\mathcal{L}$ between the full-precision and quantized LLMs.
@@ 3.1 Preserving the KV Cache of Pivot Tokens
According to Section 2.2, the attention sinks of pivot tokens are likely to deteriorate by quantization. To alleviate this issue, we propose IntactKV, a simple yet effective strategy to keep these pivot tokens intact. Specifically, as illustrated in Figure 3(a), we leverage the full-precision LLM to generate the lossless KV cache of pivot tokens, which is saved offline. The quantized LLM then loads IntactKV as the prefix to concatenate with the rest of the KV cache and continues with the regular auto-regressive decoding process. The pseudo code of the inference scheme with IntactKV is presented in Figure 3(b).
In order to study the benefits of IntactKV, we conduct a preliminary test on the mean squared error (MSE) of the attention and transformer layer output. From Figure 2, it is natural that the increasing size of IntactKV gives the monotonically decreasing MSE on both the attention and transformer layers. More importantly, it is found the pivot tokens observed in Section 2.2 (e.g., [BOS] and other delimiter tokens) give the most significant decrease on the MSE, which demonstrates the importance of their KV cache. This aligns with the observations in Figure 1 that pivot tokens exhibit outliers with extreme values and attention sinks.
@@ 3.2 IntactKV as Trainable Parameters
Since IntactKV is pre-computed and saved offline, it can be treated as extra trainable parameters aside from the LLM backbone to further boost the quantized LLMs. Despite there being no information loss at the pivot tokens, the quantization may still introduce errors to the KV cache during the decoding stage. As shown in Figure 3(a), we calibrate IntactKV to compensate for the quantization error accumulated in the following tokens. While there are various metrics to characterize the quantization discrepancy Frantar et al. (2022); Shao et al. (2024); Liu et al. (2023a), we adopt the mean squared error of the transformer layer output between the full-precision LLM and quantized LLM, a simple yet most widely used metric, i.e.,
@@
where $\Theta$ denotes the set of IntactKV, $f_{l}$ is the mapping function for the $l$-th Transformer layer, and $L$ is the number of Transformer layers in LLM. $\mathbf{x}$ is the input sequence, while $\mathbf{w},\hat{\mathbf{w}}$ are full-precision and quantized weights respectively. Note that the full-precision model is only required during the calibration process, and it can be safely discarded afterward. It is empirically found that calibration of system prompt IntactKV in SFT models generally gives more improvement than the calibration of [BOS] IntactKV in pre-trained LLMs. This matches the intuition that a larger size of IntactKV increases the potential to compensate for quantization errors.
As we focus on the post-training quantization, the training of IntactKV is highly lightweight since the only learnable parameters introduced are IntactKV, i.e., the KV cache of pivot tokens. It takes only as few as 20 epochs on a calibration set with 128 samples. Besides, training with a quantized model further lowers the memory cost. The calibration process takes about only 10 minutes for a 7B model and less than 20 minutes for a 13B model on one computing device.
@@ 3.3 Theoretical Analysis
The proof to Theorem 1 can be found in Appendix IntactKV: Improving Large Language Model Quantization by Keeping Pivot Tokens Intact. We preserve the terms w.r.t. $\Delta{\bm{K}}$ and $\Delta{\bm{V}}$ of interests, and leave the rest as constants. Note that $\Delta{\bm{K}}$ can be further separated by the pivot tokens $\Delta{\bm{K}}_{p}$ and rest tokens $\𝑝\Delta{\bm{K}}_{\textbackslash p}$, and similar notations hold for $\Delta{\bm{V}}$. Therefore, we have ${\|\Delta{\bm{K}}\|}_{2,\infty}=\max\big{(}{\|\Delta{\bm{K}}_{p}\|}_{2,\infty},{\|\Delta{\bm{K}}_{\textbackslash p}\|}_{2,\infty}\big{)},$ and ${\|\Delta{\bm{V}}\|}_{F}=\sqrt{{\|\Delta{\bm{V}}_{p}\|}_{F}^{2}+{\|\Delta{\bm{V}}_{\textbackslash p}\|}_{F}^{2}}$. With IntactKV  we have ${\|\Delta{\bm{K}}_{p}\|}_{2,\infty}=\|\Delta{\bm{V}}_{p}\|_{F}=0$ since they are generated losslessly, which decreases the upper bound of $\|\Delta{\bm{h}}\|_{2}$. Moreover, it can further reduce the bound by incorporating more pivot tokens. This also aligns with the observation in Figure 2 that a larger size of IntactKV gives a lower MSE of the attention module.
@@ 4 Experiments
FP16 7.36 6.82 6.15 5.83 7.28 6.75 5.73
RTN 9.15 7.89 6.85 6.33 8.97 7.60 6.27
    +IntactKV${}_{\textrm{[B]}}$ 8.52 7.66 6.69 6.20 8.61 7.48 6.13
GPTQ 8.59 7.49 6.73 6.29 9.58 7.43 6.33
    +IntactKV${}_{\textrm{[B]}}$ 8.30 7.42 6.62 6.23 9.27 7.36 6.28
OmniQuant 8.26 7.39 6.65 6.18 8.35 7.43 6.12
    +IntactKV${}_{\textrm{[B]}}$ 8.25 7.39 6.64 6.18 8.33 7.40 6.11
AWQ 8.26 7.38 6.59 6.16 8.31 7.32 6.05
    +IntactKV${}_{\textrm{[B]}}$ 8.12 7.36 6.54 6.12 8.18 7.29 6.04
Table 1: INT3-group128 weight-only quantization results of LLaMA and LLaMA-2 Models on C4 dataset.
@@
RTN 44.62% 51.44% 39.33% 44.56% 53.18% 61.36% 66.12% 59.05% 63.43% 67.33%
    +IntactKV${}_{\textrm{[B]}}$ 45.93% 51.89% 41.74% 46.73% 55.20% 61.94% 65.91% 61.26% 63.94% 67.95%
GPTQ 43.99% 52.95% 40.12% 47.83% 55.84% 58.61% 66.34% 59.56% 65.11% 66.66%
    +IntactKV${}_{\textrm{[B]}}$ 44.86% 52.49% 41.55% 48.53% 56.32% 59.12% 66.53% 60.46% 65.13% 67.93%
OmniQuant 46.62% 52.82% 42.95% 48.23% 55.21% 62.30% 65.58% 60.89% 64.62% 67.61%
    +IntactKV${}_{\textrm{[B]}}$ 46.27% 52.67% 43.85% 48.31% 55.51% 62.01% 65.67% 60.66% 64.89% 67.61%
AWQ 46.45% 52.92% 43.08% 48.56% 56.09% 62.18% 66.51% 60.75% 64.56% 67.67%
@@
Table 2: INT3-group128 weight-only quantization results of Vicuna models on 5-shot MMLU and 0-shot QA tasks.
@@ 4.1 Settings
We evaluate the proposed IntactKV on various sizes of open-sourced LLMs, including LLaMA (Touvron et al., 2023a) (7B-65B), LLaMA-2 (Touvron et al., 2023b) (7B-70B), Vicuna-v1.3 (Chiang et al., 2023) (7B-33B) and Vicuna-v1.5 (7B-13B). We denote models that keep intact [BOS] KV as IntactKV${}_{\textrm{[B]}}$, and models that keep intact system prompt KV as IntactKV${}_{\textrm{[P]}}$.
@@
We mainly consider weight-only quantization methods, including round-to-nearest quantization (RTN), GPTQ (Frantar et al., 2022), the state-of-the-art OmniQuant (Shao et al., 2024) and AWQ (Lin et al., 2023). For GPTQ, we use AutoGPTQ with C4 calibration set following (Frantar et al., 2022) to reproduce all results. For AWQ and OmniQuant, we use the official code or checkpoint with Pile (Gao et al., 2020) and WikiText2 (Merity et al., 2016) calibration set respectively, following (Lin et al., 2023; Shao et al., 2024). More implementation details can be found in Appendix E. We adopt asymmetric group-wise quantization with a group size of 128 and mainly focus on INT3 and INT4 quantization since INT8 is empirically lossless on various task metrics (Dettmers et al., 2022).
Our IntactKV can be readily combined with these existing weight-only quantization methods, and the experiment results are shown in Section 4.2. Moreover, aside from weight-only quantization, the proposed IntactKV can be similarly applied for KV cache quantization and extended to activation quantization, as detailed in Section 4.3 and Section 4.4. It is worth noting that the integration of IntactKV with weight-only/KV cache/activation quantization comes with no extra inference cost and works as an effective plugin to effectively boost the accuracy of quantized models.
@@
For pre-trained LLMs (i.e., LLaMA and LLaMA-2), we report the perplexity (PPL) of language generation on C4 (Raffel et al., 2020) and WikiText2 (Merity et al., 2016) dataset. For SFT models (i.e., Vicuna-v1.3 and v1.5), we conduct evaluation over a wide range of downstream tasks. We test the zero and five-shot performance on the Massively Multitask Language Understanding (MMLU) (Hendrycks et al., 2020) benchmark. Meanwhile, we also evaluate seven zero-shot commonsense QA tasks: OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), ARC-Challenge, ARC-Easy (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), and LAMBADA (Paperno et al., 2016). Additionally, we evaluate quantized Vicuna on MT-bench (Zheng et al., 2023), a high-quality dataset consisting of 80 open-ended multi-turn questions, to gauge their alignment with human preferences. The responses generated by quantized models are judged by GPT-4 with a total score of 10. More evaluation details can be found in Appendix F.
@@
Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg Hums STEM Social Others Avg
FP16 47.89% 39.96% 58.86% 57.34% 50.77% 49.78% 40.46% 60.61% 58.24% 52.10%
    +IntactKV${}_{\textrm{[B]}}$ 42.49% 35.35% 50.37% 52.44% 44.98% 44.65% 36.98% 53.04% 52.84% 46.73%
  +IntactKV${}_{\textrm{[B]}}$ 44.72% 35.42% 52.94% 52.07% 46.22% 45.61% 38.34% 55.83% 55.31% 48.53%
OmniQuant 43.51% 36.85% 52.16% 53.05% 46.18% 45.91% 37.44% 55.31% 54.94% 48.23%
    +IntactKV${}_{\textrm{[B]}}$ 44.19% 36.61% 53.33% 53.52% 46.72% 46.27% 37.54% 54.99% 54.94% 48.31%
Table 3: INT3-group128 weight-only quantization results of Vicuna-v1.3-13B on MMLU benchmarks.
#bits Method OBQA WinoGrande ARC-C ARC-E BoolQ HellaSwag LAMBADA Avg
FP16 - 45.40% 71.03% 47.70% 73.70% 82.81% 77.00% 72.91% 67.22%
w3g128 RTN 44.00% 70.96% 44.03% 67.30% 80.40% 73.33% 64.00% 63.43%
  +IntactKV${}_{\textrm{[B]}}$ 44.80% 69.93% 45.05% 68.35% 79.42% 74.81% 65.22% 63.94%
GPTQ 45.20% 69.77% 46.08% 70.33% 81.90% 74.89% 67.59% 65.11%
  +IntactKV${}_{\textrm{[B]}}$ 44.00% 70.80% 44.97% 70.75% 81.35% 75.03% 69.03% 65.13%
OmniQuant 45.20% 69.22% 45.22% 68.90% 80.95% 74.72% 68.15% 64.62%
  +IntactKV${}_{\textrm{[B]}}$ 45.40% 70.32% 45.31% 68.86% 81.28% 74.52% 68.52% 64.89%
AWQ 42.80% 68.98% 46.08% 68.98% 81.31% 74.97% 68.78% 64.56%
w4g128 RTN 45.20% 71.43% 48.04% 73.15% 82.87% 76.56% 70.62% 66.84%
  +IntactKV${}_{\textrm{[B]}}$ 44.80% 71.51% 47.44% 73.36% 82.75% 77.01% 70.99% 66.84%
GPTQ 44.60% 70.01% 47.87% 73.32% 82.23% 76.55% 71.78% 66.62%
  +IntactKV${}_{\textrm{[B]}}$ 45.00% 71.35% 46.76% 73.02% 83.33% 77.00% 71.55% 66.86%
OmniQuant 45.60% 70.56% 46.76% 73.02% 82.81% 76.74% 70.41% 66.56%
  +IntactKV${}_{\textrm{[B]}}$ 45.20% 71.43% 46.25% 72.52% 82.63% 76.90% 70.31% 66.46%
AWQ 45.20% 70.32% 47.27% 73.91% 82.81% 76.79% 71.32% 66.80%
Table 4: Weight-only quantization results of Vicuna-v1.3-13B on seven 0-shot commonsense QA tasks.
For evaluation on PPL, MMLU, and commonsense QA tasks, we adopt IntactKV${}_{\textrm{[B]}}$ that only includes [BOS] KV since the input sequence of these tasks does not use any system prompt. For evaluation of SFT models on MT-bench, we adopt IntactKV${}_{\textrm{[P]}}$ to keep an intact system prompt KV cache. The system prompt of Vicuna can be found in Appendix B. For training the cached IntactKV, we randomly sample 128 samples from ShareGPT [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered dataset as our calibration dataset, consisting of multi-turn ChatGPT (OpenAI, 2022) conversations. The layer-wise MSE loss defined in Equation 2 is calculated on the response of ChatGPT. We use AdamW optimizer with learning rate $2\times 10^{-4}$, training for 160 optimizer update steps with a gradient accumulation step of 16, i.e., 20 epochs. As mentioned in Section 3.2, training IntactKV${}_{\textrm{[B]}}$ leads to comparable performance compared with vanilla IntactKV. Instead, the calibration of IntactKV${}_{\textrm{[P]}}$ has more potential to improve quantized LLMs with longer system prompt. Thus, we primarily evaluate the IntactKV${}_{\textrm{[P]}}$ with KV cache of system prompt as trainable parameters in the following experiments. For weight and activation quantization, we further quantize IntactKV to lower bits to avoid extra inference overhead, which only incurs negligible accuracy loss. More details of activation quantization can be found in Section 4.4.